OpenStudy (anonymous):

last question: pleaseee help :/ The displacement of a mass suspended on a spring is modeled by the function y = 3sin(4)(pi)(t), where y is measured in inches and t in seconds. The amplitude, period, and frequency of the motion of the mass are given by A. amplitude = 2, period = 2, and frequency = B. amplitude = 4, period = 2, and frequency = 1 C. amplitude = 3, period = , and frequency = 4 D. amplitude = 3, period = , and frequency = 2

OpenStudy (anonymous):

amplitude = 3 w = 4pi t = 2pi/4pi = 1/2 freq = 1/t = 2

OpenStudy (anonymous):

so which is it then?

OpenStudy (anonymous):

deeeeeeeee

OpenStudy (anonymous):

hey u dint mention time period for c nd d ?